OpenStudy (anonymous):

Indefinite integral wrt x of e^x/sqrt(1-e^(2x))?

OpenStudy (anonymous):

arcsine (e^x)

OpenStudy (anonymous):

Plus a constant. Substitution used?

OpenStudy (anonymous):

\[u=e^x\] \[du=e^xdx\] gives \[\int \frac{du}{\sqrt{1-u^2}}\]

OpenStudy (anonymous):

get answer right away

OpenStudy (anonymous):

Recall derivative\[\sin^{-1} u =[1/\sqrt{1-u ^{2}}](du/dx)\]